OpenStudy (anonymous):

If you were to use the substitution method to solve the following system, choose the new system of equations that would result if x was isolated in the first equation. x + 6y + z = 2 x – 3y + 2z = –1 x + 5y + 3z = –4

OpenStudy (turingtest):

x + 6y + z = 2 x – 3y + 2z = –1 x + 5y + 3z = –4 is everything really equal to -4 like this, or are there supposed to be separate lines?